number of ways in which 9 different prizes can be given to 5 students if one particular student recieves 4 prizes abd the rest of the students can get any numbers of prizes is

Grade:11

1 Answers

Teethi
6 years ago
No.of ways to distribute 9 prizes among 4 students=9^C4Bo. of ways to distribute remaining 5 prizes among other 4 students= 4^5=2^10Hence, our final answer = 9^C4 * 2^10